OpenStudy (andertj21):

Find two consecutive odd numbers such that the sum of the first number and one-third of the second number is equal to fifty.

OpenStudy (andertj21):

35 and 37 37 and 39 49 and 51

OpenStudy (texaschic101):

2 consecutive odd numbers... x and x + 2 x + 1/3(x + 2) = 50 -- distribute x + 1/3x + 2/3 = 50 -- multiply by 3 to get rid of fractions 3x + x + 2 = 150 -- subtract 2 from both sides 3x + x = 150 - 2 4x = 148 x = 37 x + 2 = 37 + 2 = 39 check.. 37 + 1/3(39) = 50 37 + 13 = 50 50 = 50 (correct) your numbers are 37,39